OpenStudy (anonymous):

A golf ball dropped onto a hard surface rebounds exactly 1/3 of the height from which it falls. How far will it rebound after hitting the surface for the third time if it is initially dropped from height of 18m?? (please help!)

OpenStudy (anonymous):

2/3m

OpenStudy (anonymous):

if the ball drops 1/3 of 18m it implies that 1/3*18=6m now the ball then rebounds back to the surface at the same assumed 1/3 hence 1/3*6=2 and for the third and final time 1/3* 2= 2/3m